OpenStudy (anonymous):

is this correct? n+8 over 3 =15 my answer was n=14

OpenStudy (hoblos):

how did you get this answer ??

OpenStudy (anonymous):

oops i ment my answer is n=4

OpenStudy (hoblos):

is it (n+8)/3 or n +(8/3) ?

OpenStudy (anonymous):

it is n+8/3

OpenStudy (hoblos):

=> n = 15 - 8/3 = 45/3 - 8/3 = 37/3

OpenStudy (anonymous):

hoblos does it the wrong way. n + 8/ 3 =15 n + 8 = 15 x 3 n+8= 45 n= 45 - 8 n= 37

OpenStudy (anonymous):

Yeah, n=37 is the right answer !!

OpenStudy (anonymous):

Good Job !! Sparkly16!!

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

make it as a good answer

OpenStudy (anonymous):

:D

OpenStudy (phi):

Just for future reference n + 8/ 3 does not mean \[ \frac{n+8}{3} \] n+8/3 means \( n + \frac{8}{3} \) Use parentheses to write it correctly: (n+8)/3

OpenStudy (anonymous):

thanks everybody :D